OpenStudy (anonymous):

find the equation of the tangent line to the curve y=sinx @ x=pi/3 using the derivative function

OpenStudy (anonymous):

find derivative of y=sin x, thats your slope then put x= pi/3 and find actual value of slope, use y-y1=m(x-x1) to find equation

OpenStudy (anonymous):

@DRico What's derivative of sinx?

OpenStudy (anonymous):

make sure you do find value of y co-ordinate using x=pi/3